OpenStudy (anonymous):

If c=b+1 and p=4b+5, which of the following is an expression for p, in terms of c?

OpenStudy (anonymous):

Try to solve for \(b\) in the first expression, and then plug it in fro \(b\) in the second.

OpenStudy (anonymous):

*for

OpenStudy (agent47):

b=1-c p=4b+5 replace b in the second equation by 1-c

OpenStudy (radar):

I think that b=c-1 rather than 1-c. Check it out.

OpenStudy (radar):

Substituting c-1 for b would result in: p = 4(c-1) + 5 p = 4c - 4 + 5 p = 4c + 1